# Evalet: Evaluating Large Language Models by Fragmenting Outputs into Functions

### v2

@@ 7.2. Harmlessness in Conversations with Users
Ensuring that LLMs refrain from providing potentially harmful responses is a critical problem (Bai et al., 2022). To understand LLM performance in this regard, we applied our evaluation approach on the HarmfulQA dataset (Bhardwaj and Poria, 2023), which consists of questions designed to elicit harmful responses from LLMs (i.e., red teaming), along with corresponding harmless and harmful model responses. We randomly sampled 200 data points and evaluated them for Harmlessness (i.e., does the assistant engage with users safely and respectfully during interactions).
@@ 7.3. Social Intelligence in Simulated Interactions
LLMs are increasingly used as agents that simulate human users and social interactions (Park et al., 2023, 2024; Zhou et al., 2023). However, one may ask: what social behaviors should these agents simulate? To investigate this, we applied our approach on a dataset generated through the SOTOPIA environment (Wang et al., 2024), which simulates negotiations between two LLM agents role-playing characters with different social goals. We randomly sampled 200 data points and evaluated them based on Social Intelligence (i.e., does the agent effectively understand, navigate, and manage social interactions with other users or agents).
@@ 8.1. Guidelines for Integrating Fragmented and Holistic Evaluations
As revealed by our user study, both fragmented (i.e., at the fragment and function-level) and holistic (i.e., at the output-level) evaluations have distinct merits. In practice, we suggest that these types of evaluation are complementary and should be employed together. Specifically, we recommend a workflow that begins with exploration of fragmented evaluations, and transitions to holistic evaluations with the option to dive back into fragments when necessary.
Start with Fragmented Evaluation on Broad Criteria: As described by study participants, functional fragmentation can surface and reveal diverse fragment-level functions that should be considered within each criterion. Through this, one can more comprehensively identify concrete aspects that should be evaluated for a criterion, including those that were not initially considered.
Iterate and Concretize Criteria with Function Examples: By exploring fragment-level functions, practitioners can refine their criteria by deciding on what functions the evaluator should continue to surface and how these should be assessed.
Zoom Out and In: Then, as participants required in the study, practitioners should “look at the bigger picture” by exploring holistic evaluations to understand more overall qualities and their similarities. As the underlying fragment-level evaluations have been corrected and aligned, practitioners can more reliably depend on the signals provided by the holistic evaluations. At this stage, practitioners can dive deep back to the fragmented evaluations when requiring more details—alternating between levels of abstraction (Suh et al., 2023).
@@ 8.2. Calibrating Trust in LLM-as-a-Judge through Verification
Our approach, functional fragmentation, supported more calibrated and nuanced use of LLM-as-a-Judge by facilitating validation at a granular yet manageable level. By validating when the evaluator returned misaligned or inconsistent evaluations, participants calibrated their trust and selectively relied on the evaluations to guide their in-depth analysis of outputs. Despite the promise of our approach, a potential limitation is that its effectiveness depends on how reliably LLM evaluator identifies all key fragments from outputs—if fragments are surfaced, users can verify their evaluations but, if not, users cannot detect these gaps without manual review. Although our technical evaluation demonstrates strong performance, with the LLM evaluator achieving around 90% recall in identifying fragments, future work could design additional safeguards for missed fragments (e.g., separate map visualization to explore fragments that were unassessed by the evaluator).
@@ 8.3. Increasing Trend in Increasingly Longer Outputs
Recent trends in LLM advancements have focused on generating increasingly longer outputs. For example, agentic systems like Manus (AI, 2024a), Genspark (AI, 2024b), or Gemini Deep Research (Google, 2024) can create complex outcomes (e.g., multi-section reports, multi-file codebases) through multi-step workflows. Recent research has focused on further increasing output length by increasing LLM’s test-time compute (Snell et al., 2024; Guo et al., 2025; Kim et al., 2025) (i.e., training models to generate more tokens to reason more about complex tasks), or extending their context windows to receive longer inputs and generate longer outputs (AI, 2025; Hsieh et al., 2024; Ding et al., 2024). Due to their length, each part of an output can exhibit drastically different levels of quality, making it particularly difficult and challenging for practitioners to make sense of model behavior from holistic judgments. As seen in the example cases in Section 7, our approach of functional fragmentation has the potential to help users break down and interpret complex outputs in these emerging scenarios.
@@ 8.4. Qualitative and Interactive Evaluation of AI
AI/ML evaluation has mostly focused on applying quantitative metrics in benchmark datasets. This has accelerated advancements by supporting objective and concrete comparisons between models and model iterations. However, as models reach exceptionally high but similar performance on these benchmarks, users have started to qualitatively compare models based on their characteristics and behaviors, and how these fit with their own needs (Dunlap et al., 2024)—referred to as “vibe checks” (Karpathy, 2025). This raises a crucial question: “how can we help users to understand and make sense of qualitative model behaviors at scale.” To tackle this problem, our work proposes functional fragmentation to focus evaluation on the individual qualitative fragment-level functions in model outputs and, in turn, support users in sensemaking of model behaviors across outputs. However, we believe that there is still a wide design space that can be further explored by future work. In particular, due to the rich body of work in sensemaking (Chau et al., 2011; André et al., 2014; Liu et al., 2024; Pirolli and Card, 2005) and explainability (Liao et al., 2020; Lai and Tan, 2019; Kaur et al., 2022; Wang et al., 2019), we propose that the HCI community is ideally positioned to tackle this problem and to integrate itself more closely in the advancement of AI models through the development of novel interactive approaches to evaluation.
@@ 8.5. Limitations
Limitations of Functional Fragmentation: Our technical evaluation showed that our approach may fail to account for the relative importance of each fragment-level function or the holistic attributes of outputs. Future work can extend our approach to assign priority ratings to each function (e.g., manually by the practitioner or automatically suggested by the system) and generate additional functions that can represent the holistic qualities of outputs.
Dataset Scale: Our user study used datasets with hundreds of samples. However, practitioners often handle larger datasets and, while our system can support these, users may struggle to navigate and interpret the numerous fragment-level functions. Future work could introduce an additional clustering step to combine similar functions and reduce complexity.
@@ 9. Conclusion
In this work, we propose Evalet, an interactive system that instantiates our evaluation approach—functional fragmentation—by disentangling LLM outputs into fragment-level functions, rating these independently, and visualizing the landscape of these fine-grained functions. By extracting and supporting exploration of functions, Evalet supports practitioners in interpreting patterns in how LLM outputs are composed and verifying that the LLM-based evaluator is consistently evaluating aspects of interest. A within-subjects study (N=10) demonstrated that, by helping practitioners verify the LLM-based evaluations, our approach allows them to calibrate their trust and selectively rely on the evaluations to identify issues in the LLM outputs. Through illustrative case studies, we further demonstrate how Evalet effectively surfaces meaningful functions across diverse dimensions and tasks (e.g., reasoning, harmlessness, social simulations). We propose that combining both fragmented and holistic evaluation approaches can substantially enhance interpretability, comprehensiveness, and reliability when evaluating the alignment of LLMs.

### v3

@@ 7.2. Harmlessness in Conversations with Users
Ensuring that LLMs refrain from providing potentially harmful responses is a critical problem (Bai et al., 2022). To understand LLM performance in this regard, we applied our approach on 200 LLM responses to questions that purposefully elicit harmful behaviors (Bhardwaj and Poria, 2023). We assess these on (i.e., the assistant engages safely and respectfully).
@@ 7.3. Social Intelligence in Simulated Interactions
LLMs are increasingly used as agents that simulate human users and social interactions (Park et al., 2023, 2024; Zhou et al., 2023). However, one may ask: what social behaviors should these agents simulate? To investigate this, we applied our approach on 200 simulated dialogues between two role-playing LLM agents based on (i.e., the agent effectively understands, navigates, and manages social interactions).
@@ 8.1. Guidelines for Integrating Fragmented and Holistic Evaluations
As revealed by our user study, both fragmented (i.e., at the fragment and function-level) and holistic (i.e., at the output-level) evaluations have distinct merits. In practice, we suggest that these types of evaluation are complementary and practitioners should employ them together through a layered workflow:
Start with Fragmented Evaluation on Broad Criteria: As described by study participants, functional fragmentation can surface diverse fragment-level functions that should be considered within each criterion. Through this, practitioners can more comprehensively identify concrete aspects to evaluate for each criterion, including those that were not initially considered.
Iterate and Concretize Criteria with Function Examples: By exploring fragment-level functions, practitioners can refine their criteria by deciding on what functions the evaluator should continue to surface and how these should be assessed. Practitioners should iteratively refine their criteria and example sets while re-evaluating the outputs until most misalignments disappear and the function clusters stabilize (i.e., similar clusters are presented each evaluation).
Zoom Out and In: Then, practitioners should zoom out to inspect the whole outputs, and their holistic justifications and scores to understand the overall qualities and similarities of outputs. As the underlying fragment-level evaluations have been corrected and aligned, practitioners can more reliably depend on the signals from the holistic scores. At this stage, practitioners can dive deep back to the fragmented evaluations when they need more details (e.g., identify concrete issues in low scoring outputs or the patterns of those with mixed scores). This allows users to flexibly alternate between levels of abstraction (Suh et al., 2023).
To fully support this integrated workflow, Evalet must also capture the holistic attributes of outputs (e.g., tone, structure). While the system does not currently support this, this can be easily addressed by introducing an additional LLM module after the functional fragmentation step that extracts holistic attributes from the output that are related to each criterion. These attributes can be represented as sentence-level descriptions (e.g., ”comedic tone”)—the same format as the function labels of fragments. Then, these holistic attribute labels can also be compared and visualized with the other fragment-level functions, and also factored into the holistic scores.
@@ 8.2. Calibrating Trust in LLM-as-a-Judge through Verification
Our approach, functional fragmentation, supported more calibrated and nuanced use of LLM-as-a-Judge by facilitating validation at a granular yet manageable level. In our study, this calibrated trust allowed participants to selectively rely on the evaluations to scaffold more in-depth analysis of outputs. Despite the promise of our approach, a potential limitation is that its effectiveness depends on how reliably the LLM evaluator identifies all key fragments from outputs—if fragments are surfaced, users can verify their evaluations but, if not, users cannot detect these gaps without manual review. Although our technical evaluation demonstrates strong performance, with the LLM evaluator achieving around 90% recall in identifying fragments, future work could design additional safeguards for missed fragments. For example, Evalet could integrate a separate map visualization that embeds fragments that were not assessed by the evaluator for any criterion to support navigation of missed fragments.
@@ 8.3. Increasing Trend in Increasingly Longer Outputs
Recent trends in LLM advancements have focused on generating increasingly longer outputs. For example, agentic systems like Manus (Meta, 2026), Genspark (AI, 2024), or Gemini Deep Research (Google, 2024) can carry out multi-step workflows to create complex outcomes (e.g., multi-section reports, multi-file codebases). Recent research looks to further increase output length by increasing LLM’s test-time compute (Snell et al., 2024; Guo et al., 2025; Kim et al., 2025) (i.e., training models to generate more tokens at once), or extending their context windows to handle longer inputs and outputs (AI, 2025; Hsieh et al., 2024; Ding et al., 2024). In longer outputs, each part of the output can exhibit drastically different levels of quality, making it particularly difficult and challenging for practitioners to make sense of model behavior from holistic judgments. We propose that functional fragmentation can help practitioners break down and interpret complex and lengthy outputs. For example, in Appendix E.4, we present an example of applying our approach to the traces from computer use agents, which surfaced meaningful behaviors such as agents using keyboard shortcuts to complete tasks more efficiently or instances where the agent performed inefficient redundant actions. Our work presents preliminary evidence of the potential of our approach in these emergent scenarios and suggests that future work can further explore this potential.
@@ 8.4. Qualitative and Interactive Evaluation of AI
AI/ML evaluation has mostly focused on applying quantitative metrics in benchmark datasets. This accelerated advancements by supporting objective and concrete comparisons between models and model iterations. However, as models reach exceptionally high but similar performance on these benchmarks, users have started to qualitatively compare models based on their behaviors and how these fit with their own needs (Dunlap et al., 2024)—referred to as “vibe checks” (Karpathy, 2025). This raises a crucial question: “how can we help users to understand and make sense of qualitative model behaviors at scale?” To tackle this problem, our work proposes functional fragmentation to focus evaluation on the individual qualitative fragment-level functions in model outputs and, in turn, support users in sensemaking of model behaviors across outputs. While benchmarks serve to monitor progress in models’ fundamental capabilities, qualitative and ad-hoc evaluation methods can complement them by characterizing model behaviors. Furthermore, interactive and qualitative evaluation methods, like functional fragmentation, are critical in tasks and domains where benchmarks do not exist. Given its rich body of work in sensemaking (Chau et al., 2011; André et al., 2014; Liu et al., 2024; Pirolli and Card, 2005) and explainability (Liao et al., 2020; Lai and Tan, 2019; Kaur et al., 2022; Wang et al., 2019), we propose that the HCI community is ideally positioned to tackle this challenge and integrate itself more closely in the advancement of AI models by developing novel interactive evaluation methods.
@@ 8.5. Limitations
Function Priority: Our technical evaluation showed that our approach does not account for the relative importance of each fragment-level function. Future work could add priority ratings to functions (e.g., manually by the user or suggested by the system).
Controllability: Evalet supports control of the functional fragmentation process by allowing users to exclude or re-rate fragment-level functions. We initially included or considered additional controls (e.g., function relabeling, cluster editing, fragment-size editing), but pilot users found that these added cognitive overhead without clear benefit. Due to the time constraints in the user study, we opted for not including these features in the system used during the study. However, as seen from the findings, some participants reported needing more control (e.g., fragment size), suggesting that this feature is valuable in certain workflows or to certain users. In the open-source release of Evalet, we plan to introduce these controls as advanced settings for users.
Dataset Scale: Our user study and examples used datasets with 100–200 samples, whereas practitioners often handle larger datasets. Although Evalet can support these, the linear growth in fragment-level functions and clusters may hinder users’ ability to navigate the space and decide where to start their analysis. One way to address this is to apply our clustering technique recursively to create more cluster levels, which can help summarize and decompose the vast space of functions for users. Evalet can also incorporate proactive agentic guidance (Pu et al., 2025; Prasongpongchai et al., 2025; Chen et al., 2025) to identify and highlight significant clusters or functions that can help kickstart users’ analysis.
@@ 9. Conclusion
In this work, we propose Evalet, an interactive system that instantiates functional fragmentation: disentangling LLM outputs into fragment-level functions and visualizing the landscape of these functions across outputs. By extracting and supporting exploration of functions, Evalet helps practitioners interpret patterns in how LLM outputs are composed and verifying that the LLM-based evaluator is consistently evaluating these aspects. A within-subjects study (N=10) demonstrated that our approach helps practitioners verify LLM evaluations and, in turn, calibrate their trust in these evaluations. Through illustrative case studies, we further demonstrate how Evalet effectively surfaces meaningful functions across diverse dimensions and tasks (e.g., reasoning, harmlessness, social simulations).
Acknowledgements.
We thank all of our participants for engaging positively in our user study. We would also like to thank the reviewers for their thoughtful feedback and comments. Finally, we would like to thank Juhoon Lee and members of KIXLAB for their help in polishing this manuscript. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (Ministry of Science and ICT) (No.RS-2025-00557726). This work was also supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00443251, Accurate and Safe Multimodal, Multilingual Personalized AI Tutors).References
- Genspark: agents that write code and explain it. Note: https://genspark.ai/Accessed: 2025-04-10 Cited by: §8.3.
- The llama 4 herd: the beginning of a new era of natively multimodal ai innovation. Note: https://ai.meta.com/blog/llama-4-multimodal-intelligence/Accessed: 2026-02-02 Cited by: §8.3.
- Crowd synthesis: extracting categories and clusters from complex data. In Proceedings of the 17th ACM conference on Computer supported cooperative work & social computing, pp. 989–998. Cited by: §8.4.
- Claude 3.7 sonnet and claude code. Note: Accessed: 2025-03-19 External Links: Link Cited by: §4.3.
- Chainforge: a visual toolkit for prompt engineering and llm hypothesis testing. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, pp. 1–18. Cited by: §1, §2.2.
- Aligning human and llm judgments: insights from evalassist on task-specific evaluations and ai-assisted assessment strategy preferences. arXiv preprint arXiv:2410.00873. Cited by: §2.2.
- Measuring implicit bias in explicitly unbiased large language models. External Links: 2402.04105, Link Cited by: §8.2.
- Constitutional ai: harmlessness from ai feedback. arXiv preprint arXiv:2212.08073. Cited by: §7.2.
- Red-teaming large language models using chain of utterances for safety-alignment. External Links: 2308.09662 Cited by: §E.2, §7.2.
- Language models are few-shot learners. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33, pp. 1877–1901. External Links: Link Cited by: §4.1.5.
- Zeno: an interactive framework for behavioral evaluation of machine learning. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI ’23, New York, NY, USA. External Links: ISBN 9781450394215, Link, Document Cited by: §2.1, §3.3.
- What did my ai learn? how data scientists make sense of model behavior. ACM Transactions on Computer-Human Interaction 30 (1), pp. 1–27. Cited by: §1.
- Booookscore: a systematic exploration of book-length summarization in the era of llms. arXiv preprint arXiv:2310.00785. Cited by: §2.2.
- Apolo: making sense of large network data by combining rich user interaction and machine learning. In Proceedings of the SIGCHI conference on human factors in computing systems, pp. 167–176. Cited by: §8.4.
- Need help? designing proactive ai assistants for programming. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, pp. 1–18. Cited by: 4th item.
- TaleBrush: sketching stories with generative pretrained language models. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems, pp. 1–19. Cited by: §1.
- Longrope: extending llm context window beyond 2 million tokens. arXiv preprint arXiv:2402.13753. Cited by: §8.3.
- Is gpt-3 text indistinguishable from human text? scarecrow: a framework for scrutinizing machine text. arXiv preprint arXiv:2107.01294. Cited by: §2.2.
- VibeCheck: discover and quantify qualitative differences in large language models. arXiv preprint arXiv:2410.12851. Cited by: §3, §8.4.
- Gptscore: evaluate as you desire. arXiv preprint arXiv:2302.04166. Cited by: §1.
- MetricMate: an interactive tool for generating evaluation criteria for llm-as-a-judge workflow. Cited by: §1, §2.2.
- Structure-mapping: a theoretical framework for analogy. Cognitive science 7 (2), pp. 155–170. Cited by: §2.3.
- Supporting sensemaking of large language model outputs at scale. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, pp. 1–21. Cited by: §2.3, §3.
- OverCode: visualizing variation in student solutions to programming problems at scale. ACM Transactions on Computer-Human Interaction (TOCHI) 22 (2), pp. 1–35. Cited by: §2.3.
- Gemini deep research - your personal research assistant. Note: https://gemini.google/overview/deep-research/Accessed: 2025-04-10 Cited by: §8.3.
- AbstractExplorer: leveraging structure-mapping theory to enhance comparative close reading at scale. In Proceedings of the 38th Annual ACM Symposium on User Interface Software and Technology, pp. 1–25. Cited by: §2.3.
- Deepseek-r1: incentivizing reasoning capability in llms via reinforcement learning. arXiv preprint arXiv:2501.12948. Cited by: §E.1, §7.1, §8.3.
- RULER: what’s the real context size of your long-context language models?. arXiv preprint arXiv:2404.06654. Cited by: §8.3.
- Openai o1 system card. arXiv preprint arXiv:2412.16720. Cited by: §1, §7.1.
- Graphologue: exploring large language model responses with interactive diagrams. In Proceedings of the 36th annual ACM symposium on user interface software and technology, pp. 1–20. Cited by: §2.3.
- LLM comparator: interactive analysis of side-by-side evaluation of large language models. IEEE Transactions on Visualization and Computer Graphics. Cited by: §2.2.
- My reaction is that there is an evaluation crisis. i don’t really know what metrics to look at right now. […] in absence of great comprehensive evals i tried to turn to vibe checks instead, but i now fear they are misleading and there is too much opportunity for confirmation bias, too low sample size, etc., it’s just not great. tldr my reaction is i don’t really know how good these models are right now.. Note: https://x.com/karpathy/status/1896266683301659068Accessed: 2025-04-01 Cited by: §8.4.
- Sensible ai: re-imagining interpretability and explainability using sensemaking theory. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency, pp. 702–714. Cited by: §8.4.
- Topiclens: efficient multi-level visual topic exploration of large-scale document collections. IEEE transactions on visualization and computer graphics 23 (1), pp. 151–160. Cited by: §2.3.
- Prometheus: inducing fine-grained evaluation capability in language models. In The Twelfth International Conference on Learning Representations, Cited by: §1, §3.2.
- The biggen bench: a principled benchmark for fine-grained evaluation of language models with language models. arXiv preprint arXiv:2406.05761. Cited by: §3.2.
- Scaling evaluation-time compute with reasoning models as process evaluators. arXiv preprint arXiv:2503.19877. Cited by: §8.3.
- Cells, generators, and lenses: design framework for object-oriented interaction with large language models. In The 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23), October 29-November 1, 2023, San Francisco, CA, USA, UIST ’23, New York, NY, USA. External Links: ISBN 9798400701320, Link, Document Cited by: §2.2, §6.1.3.
- Evallm: interactive evaluation of large language model prompts on user-defined criteria. In Proceedings of the CHI Conference on Human Factors in Computing Systems, pp. 1–21. Cited by: §1, §1, §2.2, §3.1, §3, 2nd item, §6.1.2.
- Human-ai collaboration via conditional delegation: a case study of content moderation. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems, pp. 1–18. Cited by: §2.1.
- On human predictions with explanations and predictions of machine learning models: a case study on deception detection. In Proceedings of the conference on fairness, accountability, and transparency, pp. 29–38. Cited by: §8.4.
- AI policy projector: grounding llm policy design in iterative mapmaking. arXiv preprint arXiv:2409.18203. Cited by: §2.3.
- Concept induction: analyzing unstructured text with high-level concepts using lloom. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, pp. 1–28. Cited by: §2.3, §4.2.2.
- Rewardbench: evaluating reward models for language modeling. arXiv preprint arXiv:2403.13787. Cited by: §B.2.1, §5.2.1.
- Questioning the ai: informing design practices for explainable ai user experiences. In Proceedings of the 2020 CHI conference on human factors in computing systems, pp. 1–15. Cited by: §2.1, §8.4.
- Human-centered explainable ai (xai): from algorithms to user experiences. arXiv preprint arXiv:2110.10790. Cited by: §2.1.
- Wildbench: benchmarking llms with challenging tasks from real users in the wild. arXiv preprint arXiv:2406.04770. Cited by: §3.2.
- “What it wants me to say”: bridging the abstraction gap between end-user programmers and code-generating large language models. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI ’23, New York, NY, USA. External Links: ISBN 9781450394215, Link, Document Cited by: §2.2.
- Selenite: scaffolding online sensemaking with comprehensive overviews elicited from large language models. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, pp. 1–26. Cited by: §8.4.
- G-eval: nlg evaluation using gpt-4 with better human alignment. arXiv preprint arXiv:2303.16634. Cited by: §1.
- Least squares quantization in pcm. IEEE transactions on information theory 28 (2), pp. 129–137. Cited by: item 3.
- The ai scientist: towards fully automated open-ended scientific discovery. arXiv preprint arXiv:2408.06292. Cited by: §1.
- A unified approach to interpreting model predictions. Advances in neural information processing systems 30. Cited by: §2.1.
- Some methods for classification and analysis of multivariate observations. In Proceedings of the Fifth Berkeley Symposium on Mathematical Statistics and Probability, Volume 1: Statistics, Vol. 5, pp. 281–298. Cited by: item 3.
- Hdbscan: hierarchical density based clustering.. J. Open Source Softw. 2 (11), pp. 205. Cited by: item 2.
- Umap: uniform manifold approximation and projection for dimension reduction. arXiv preprint arXiv:1802.03426. Cited by: item 1.
- Manus: hands on ai. Note: https://manus.imAccessed: 2026-02-02 Cited by: §8.3.
- Factscore: fine-grained atomic evaluation of factual precision in long form text generation. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pp. 12076–12100. Cited by: §2.2.
- PromptAid: prompt exploration, perturbation, testing and iteration using visual analytics for large language models. arXiv preprint arXiv:2304.01964. Cited by: §2.2.
- Qualeval: qualitative evaluation for model improvement. arXiv preprint arXiv:2311.02807. Cited by: §2.1.
- Evaluating content selection in summarization: the pyramid method. In Proceedings of the human language technology conference of the north american chapter of the association for computational linguistics: Hlt-naacl 2004, pp. 145–152. Cited by: §2.2.
- CoNotate: suggesting queries based on notes promotes knowledge discovery. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems, pp. 1–14. Cited by: §2.3.
- InterWeave: presenting search suggestions in context scaffolds information search and synthesis. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology, pp. 1–16. Cited by: §2.3.
- Generative agents: interactive simulacra of human behavior. In Proceedings of the 36th annual acm symposium on user interface software and technology, pp. 1–22. Cited by: §7.3.
- Generative agent simulations of 1,000 people. arXiv preprint arXiv:2411.10109. Cited by: §7.3.
- Performance assessment: designing appropriate performance tasks and scoring rubrics.. Cited by: §3.2, §3.
- The sensemaking process and leverage points for analyst technology as identified through cognitive task analysis. In Proceedings of international conference on intelligence analysis, Vol. 5, pp. 2–4. Cited by: §2.3, §8.4.
- Talk to the hand: an llm-powered chatbot with visual pointer as proactive companion for on-screen tasks. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, pp. 1–16. Cited by: 4th item.
- Assistance or disruption? exploring and evaluating the design and trade-offs of proactive ai programming support. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, pp. 1–21. Cited by: 4th item.
- Forsense: accelerating online research through sensemaking integration and machine research support. In Proceedings of the 26th International Conference on Intelligent User Interfaces, pp. 608–618. Cited by: §2.3.
- Adaptive testing and debugging of nlp models. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pp. 3253–3267. Cited by: §2.1.
- ”Why should i trust you?” explaining the predictions of any classifier. In Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining, pp. 1135–1144. Cited by: §2.1.
- Beyond accuracy: behavioral testing of nlp models with checklist. arXiv preprint arXiv:2005.04118. Cited by: §1, §2.1, §3.2, §3.
- Angler: helping machine translation practitioners prioritize model improvements. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI ’23, New York, NY, USA. External Links: ISBN 9781450394215, Link, Document Cited by: §2.1.
- Lmunit: fine-grained evaluation with natural language unit tests. arXiv preprint arXiv:2412.13091. Cited by: §3.2.
- Who validates the validators? aligning llm-assisted evaluation of llm outputs with human preferences. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology, pp. 1–14. Cited by: §1, §2.2.
- Towards bidirectional human-ai alignment: a systematic review for clarifications, framework, and future directions. arXiv preprint arXiv:2406.09264. Cited by: §1.
- Divisi: interactive search and visualization for scalable exploratory subgroup analysis. arXiv preprint arXiv:2502.10537. Cited by: §2.1, §3.3.
- Scaling llm test-time compute optimally can be more effective than scaling model parameters. arXiv preprint arXiv:2408.03314. Cited by: §1, §8.3.
- PaperBench: evaluating ai’s ability to replicate ai research. arXiv preprint arXiv:2504.01848. Cited by: §1.
- Interactive and visual prompt engineering for ad-hoc task adaptation with large language models. IEEE Transactions on Visualization and Computer Graphics 29 (1), pp. 1146–1156. External Links: Document Cited by: §2.2.
- Luminate: structured generation and exploration of design space with large language models for human-ai co-creation. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, pp. 1–26. Cited by: §2.3, §6.1.3.
- Sensecape: enabling multilevel exploration and sensemaking with large language models. In Proceedings of the 36th annual ACM symposium on user interface software and technology, pp. 1–18. Cited by: §2.3, item 3.
- Comparing criteria development across domain experts, lay users, and models in large language model evaluation. arXiv preprint arXiv:2410.02054. Cited by: §1.
- Clio: privacy-preserving insights into real-world ai use. arXiv preprint arXiv:2412.13678. Cited by: §4.2.2.
- Open Thoughts. Note: https://open-thoughts.ai Cited by: §E.1, §7.1.
- A general inductive approach for analyzing qualitative evaluation data. American journal of evaluation 27 (2), pp. 237–246. Cited by: §3.1, §3.
- Designing theory-driven user-centric explainable ai. In Proceedings of the 2019 CHI conference on human factors in computing systems, pp. 1–15. Cited by: §8.4.
- SOTOPIA-: interactive learning of socially intelligent language agents. arXiv preprint arXiv:2403.08715. Cited by: §E.3.
- Opencua: open foundations for computer-use agents. arXiv preprint arXiv:2508.09123. Cited by: §E.4.
- Wizmap: scalable interactive visualization for exploring large machine learning embeddings. arXiv preprint arXiv:2306.09328. Cited by: §2.3.
- The what-if tool: interactive probing of machine learning models. IEEE Transactions on Visualization and Computer Graphics 26 (1), pp. 56–65. External Links: Document Cited by: §2.1.
- ScatterShot: interactive in-context example curation for text transformation. In Proceedings of the 28th International Conference on Intelligent User Interfaces, IUI ’23, New York, NY, USA, pp. 353–367. External Links: ISBN 9798400701061, Link, Document Cited by: §2.2.
- Polyjuice: generating counterfactuals for explaining, evaluating, and improving models. arXiv preprint arXiv:2101.00288. Cited by: §2.1.
- Errudite: scalable, reproducible, and testable error analysis. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Florence, Italy, pp. 747–763. External Links: Link, Document Cited by: §2.1, §3.3.
- AI chains: transparent and controllable human-ai interaction by chaining large language model prompts. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems, CHI ’22, New York, NY, USA. External Links: ISBN 9781450391573, Link, Document Cited by: §2.2.
- Fine-grained human feedback gives better rewards for language model training. arXiv preprint arXiv:2306.01693. Cited by: §B.1.1, §5.1.1.
- Review spotlight: a user interface for summarizing user-generated reviews using adjective-noun word pairs. In Proceedings of the SIGCHI conference on human factors in computing systems, pp. 1541–1550. Cited by: §2.3.
- Flask: fine-grained language model evaluation based on alignment skill sets. arXiv preprint arXiv:2307.10928. Cited by: §3.2, §3.
- Wordcraft: story writing with large language models. In Proceedings of the 27th International Conference on Intelligent User Interfaces, pp. 841–852. Cited by: §1, §6.1.3.
- Herding ai cats: lessons from designing a chatbot by prompting gpt-3. In Proceedings of the 2023 ACM Designing Interactive Systems Conference, DIS ’23, New York, NY, USA, pp. 2206–2220. External Links: ISBN 9781450398930, Link, Document Cited by: §2.2.
- EvalTree: profiling language model weaknesses via hierarchical capability trees. arXiv preprint arXiv:2503.08893. Cited by: §2.1.
- ChainBuddy: an ai agent system for generating llm pipelines. arXiv preprint arXiv:2409.13588. Cited by: §2.2.
- Judging llm-as-a-judge with mt-bench and chatbot arena. External Links: 2306.05685 Cited by: §1, §2.2, §3, §6.1.2.
- Towards a unified multi-dimensional evaluator for text generation. arXiv preprint arXiv:2210.07197. Cited by: §1.
- Sotopia: interactive evaluation for social intelligence in language agents. arXiv preprint arXiv:2310.11667. Cited by: §7.3.